Journalist Report – December 19th

[Sol 1]
And we have landing! This morning, we arrived on Mars as expected and we began our mission, which will last until the 30th of December with the arrival of Crew 186. But right now, we are alone in MDRS and our crew is in a great shape and a dynamic mentality.
This morning, we began the day with an engineering check to verify that the generator, the rovers, the water system… were in a good condition. It was the first time with an EVA suit for Arno and David, guided by our experienced Commander Ilaria. The station is apparently doing well and we hope that it will keep working that way until the end of the mission.
After a common lunch, we did an important meeting to set up ou schedule for the upcoming days. Then, David went to the Greehab to move his experiences forward, while our spacesuit engineers repaired two of the EVA suits that had jammed ventilators. Right now, we have six functionnal suits for five people, which is promising!
After dinner, we will do a mindfulness training and meditation and we will work on the different experiments that we will implement and run. We also hope to be able to watch a (space) movie together.
Greetings from Mars,
Thibault, ExO and spacesuit engineer for the Crew 185

Sol Summary – December 19th

Crew 185 Daily Summary Report  19/12/2017
MDRS Daily Summary Report for Sol 1
First Day of Sim
Mission Status:  Things are running smoothly.
Sol Activity Summary: We started our simulation this morning. We began our day by eating breakfast as a team and followed our meal with a mindfulness meditation session. After that, three of us put on our simulation space suits for the first time and left the hab to do an engineering check. Two crew members stayed behind for communication. When the crew returned to the hab, we ate lunch, assigned daily tasks for each crew member, and had time to work on individual projects.
Look Ahead Plan: Tomorrow, we will have our first EVA and test the X-1 space suit for the first time.
Anomalies in work:  The simulation space suit helmet fogs up easily and can drastically limit your field of vision.
David Murray crew greenhab officer and biologist
Weather: Sunny, warm
Crew Physical Status: Healthy
EVA: No EVA today
Reports to be file: EVA request
Support Requested: None

Operations Report – December 18th

Crew 185
Operations Report Monday 12/18/2017
SOL: -1
Name of person filing report: A.Passeron
Non-nominal systems:
– Heater :
On Monday morning, Thibault and I removed the air filter. It had not been changed since April 2016. At 10:00am, Ilaria and David went to city to buy a battery for the generator and a pack of air filters. We replaced the filter at 3:00pm. The heater is now working well.
-Generator :
Shannon helped us to change the battery at 3:00pm. We then drained oil from the radiator by pouring clean cooling liquid. The oil of the generator was also very dirty. It was so sticky that we had to start the generator for a few minutes to heat the oil and thin it. We thus were able to remove the oil filter, but only a fraction of dirty oil came out of the generator. We tried to remove as much dirty oil as possible by filling the motor with clean oil, but a valve seemed to block the oil circulation. After pouring as much clean oil as we could, we replaced the oil filter. The generator was finally turned on at 5:00pm.
Generator (hours run):
Generator turned on at 5:00pm.
Solar— SOC @ 6:00pm : 98%
Diesel – 80%
Propane – 60 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 Gallons (Full)
Water (trailer) – 150 gallons
Water (static) – 300 gallons
Trailer to Static Pump used – Yes
Static to Loft Pump used – Yes
Water Meter: 45 gallons
Toilet tank emptied: No
ATV’s Used: No
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Opportunity rover used:  No
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: All Nominal
Summary of Hab operations: All nominal
Summary of GreenHab operations: All nominal
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

Sol Summary – December 18th

Today has been a busy day for the Crew 185. As it was our last day before entering simulation, we needed to prepare the hab in the best possible way and to fix as many things as possible before closing our doors.

We decided to work in two groups. Ilaria and David took the car to buy a new battery for the generator, soil for the greenhab and food. Arno, Thibault and John set the laboratory with the 3D printer, the medical supplies, the X-1 spacesuit prototype, etc…

This afternoon, we repaired the generator by replacing the battery and changing a part of the cooling liquid and the oil. It is working better at the moment when I write these lines! Hopefully, it will keep being in a good shape.

We decided that we would begin the simulation tomorrow morning when waking up. Mars, here we come!

The Crew 185

Operations Report – December 17th

Crew 185
Operations Report Sunday 12/17/2017
SOL: SOL -2
Name of person filing report: A.Passeron
Non-nominal systems:
– Heater :
The heater was not working properly, even after reseting it. Shannon adviced us to change the air filter.
– Generator :
The generator had a failure at 8:00pm, Fault message reading “Low battery”. Jump-start from Shannon’s Shuttle battery did not solve the issue. Generator had a second failure at 9:00pm, so we definitely turned it off for the night.
Generator (hours run):
Generator turned off at 9:00am.
Generator turned on at 5:00pm.
Generator had a failure at 8:00pm.
Jump-start done at 8:00pm.
Generator turned off at 9:00pm after a second failure.
Solar— SOC @ 5:00pm : 95%
Diesel – 80%
Propane – 62 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 Gallons (Full)
Water (trailer) – 300 gallons
Water (static) – 230 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: Yes
ATV’s Used: Yes
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:40 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  Yes
Hours: 00:40 Hours
Beginning charge: 100%
Ending charge: 90%
Currently charging: Yes
Opportunity rover used:  Yes
Hours: 00:40 Hours
Beginning charge: 100%
Ending charge: 88%
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: All Nominal
Summary of Hab operations: All nominal
Summary of GreenHab operations: Greenhab report is nominal and cleaned!
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

GreenHab Report – December 17th

Greenhab Report
David Murray
December 17th, 2017
Environmental control: (Choose which is appropriate and explain further if needed)
Ambient
Ambient with window/door open from 10:30 – ? (Someone closed door without notifying me)
Working Hour: 18:30
Outside temp during working hours: 1 C
Inside temperature high: 31 C
Inside temperature low: 15 C
Inside humidity: 16%
Heating
Functioning nominally
Turns on at 16 C and off at 20 C
Cooling
Not functional
Shade Cloth On/Off
On
Average temperature:
N/A
Changes to Crops
Bean plants are no longer drooping
One bean plant has 9 flowers open or partially open
Two bean plants have one flower open or partially open
One bean plant has two flowers open or partially open
Dill weed has sprouted
Morning Research Observations
N/A
Daily Water Usage for Crops
5 gallons
Time(s) of Watering Crops
Around 13:00 and 18:30
Aquaponics
N/A
Narrative
Took inventory of all identifiable living plants in the greenhab.
Cucumber- 23
Tomato- 101
Beans- 31
Melon- 21
Pepper- 25
Many tomatoes need to be transplanted into larger pots
Support/Supplies Needed
N/A

Operations Report – December 16th

Crew 185
Operations Report Saturday 12/16/2017
SOL: -3
Name of person filing report: APasseron
Non-nominal systems: None
Notes on non-nominal systems:
Generator (hours run): Generator is running and nominal.
Solar— SOC @ 9:00pm : 100%
Diesel – 85%
Propane – 62 percent volume
Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 Gallons (Full)
Water (trailer) – 400 gallons
Water (static) – 250 gallons
Trailer to Static Pump used – No
Static to Loft Pump used – Yes
Water Meter: 50 gallons
Toilet tank emptied: No
ATV’s Used: None
Oil Added? No
ATV Fuel Used:
# Hours the ATVs were used today: 00:00 hours
Notes on ATVs: ATVs were nominal.
Deimos rover used: No
Hours: 00:00 hours
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director only.
Hours: 00:00 Hours
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used:  Yes
Hours: 00:30 Hours
Beginning charge: 100%
Ending charge: 80%
Currently charging: Yes
Opportunity rover used:  Yes
Hours: 00:30 Hours
Beginning charge: 100%
Ending charge: 80%
Currently charging: Yes
Curiosity rover used:  No
Hours:
Beginning charge:
Ending charge:
Currently charging:
HabCar used and why, where?  No
Summary of internet: All Nominal
Summary of suits and radios: All Nominal
Summary of Hab operations: All nominal
Summary of GreenHab operations: Greenhab report is nominal and cleaned!
Summary of ScienceDome operations: All nominal
Summary of RAMM operations: Not Operational
Summary of health and safety issues: Crew is Healthy
Questions, concerns and requests to Mission Support:

GreenHab Report – December 16th

GreenHab Report

Written by David Murray

Supervised by Trisha Randazzo

December 16, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Outside temperature during work hours:  -4.5 C

·         At 20:25 inside temperature was 15 C.

·         Inside temperature high 31

·         Inside temperature low 15

·         Inside humidity: 16%

Heating

·         Functioning nominally

·         Turns on at 16 C stops at 20 C

Cooling

·         Not available

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         Two of the bean plants are still drooping

·         The sprouting bean plants now has 4 flowers

Daily water usage for crops:

·         Used more water to accommodate the drier planter box.

Time(s) of watering for crops:

·         20:25

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·         4 Gallons

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         Completed crew handover activities for GreenHab. David is now well trained as the new GreenHab Officer.

Support/supplies needed:

·         Soil and a fire extinguisher

Astronomy Report – December 15th

Astronomy Report

Name:  Thomas Horn    Crew: 184
Date: 7/15/17

Sky Conditions: Clear

Wind Conditions: Light

Observation Start Time: 2:00 pm

Observation End Time: 2:30 pm

Summary: No observing, shutdown Observatory and returned Laptop to HAB

Objects Viewed: None

Problems Encountered: None

Sol Summary – December 15th

MDRS Sol Summary Report for Sol 14

Summary Title: Signing off!

Mission Status:

It’s been a please working with all of you, and thanks for your help.

It was a fitting send off to our team that our day started with a
generator failure. Upon our morning rounds after waking up it was
noticed the Generator had failed with an ‘Underspeed’ Fault. We were
happy to see that SOC was holding steady at 100% (thank you solar
arrays!) but this was a worrying sign, especially as we didn’t want to
hand over a crippled ship to the oncoming crew.

Investigation into the generator showed that the loose connections
that were repaired previously were still holding strong, so there was
no immediate cause of the generator fault. After attempting to
restart the generator a low battery voltage was observed (reading
7.6volts, supposed to be 9.6 volts). It appears that after the
generator failed it drained the battery down displaying the fault
message on its status plate. After a jumpstart from our friendly
neighbors we were able to get the generator back up and running. This
issue is likely to re-occur as we are still not sure of the original
cause. Our next step will be to Remove and Replace the generator
batteries, fingers crossed this fixes the problem. Best of luck to
future crews in continuing to troubleshoot this issue.

Besides that morning excitement the day was one of cleaning and
packing, with a short break in the afternoon to do a short EVA around
the HAB for some last minute filming to get the shots we’ve missed so
far this mission. We hope to leave this place better than we found
it.

This report comes to you with the assistance of Mission Support as our
Internet allowance has run out. Sometime today the same laptop that
killed our bandwidth before had WiFi turned On, after which it sucked
our bandwidth dry again. Still don’t know why this particular laptop
uses bandwidth at such a prodigious rate, but we’ve turned it off
again and comms should be returned to normal.

Hope to work with you all again in the future.

Ad Astra Per Ardua,

Crew 184.

Sol Activity Summary:

1. Filming EVA

2. Cleaning / Packing

4. Reports and general HAB maintenance / checks

6. Generator Failure and Jumpstart

Look Ahead Plan:

1. Crew Handover Training

Anomalies in work:

1. HAB Leak from under EVA door continuing (per Mission Control
leak is expected and has been present for months, presumed to be from
Kitchen piping)

2. Generator Underspeed Fault

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.

EVA: None today.

Reports to be filed:

Greenhab

Operations

Journalist

Astronomy

Sol Summary

Mission Summary

Support Requested:

1. None